OpenStudy (anonymous):

When one U.S. Dollar is worth one point four eight Australian Dollars, which of the following could be used as a conversion factor? A. 250/37 B. 1 12/25 C. 1 29/ 50 D. 2 13/25

OpenStudy (anonymous):

Which fraction gives you 1.48 ?

OpenStudy (ajprincess):

Hint:Try converting 1.48 as mixed fraction. Does that help? @Gigglez

OpenStudy (anonymous):

o.o i hate fractions...

OpenStudy (ajprincess):

\[1.48=\frac{148}{100}\]Can u do it nw?

OpenStudy (ajprincess):

@Gigglez

OpenStudy (anonymous):

Ughhhhh rawrrarrrrrrrrrrrrrrrr

OpenStudy (ajprincess):

@Gigglez may be u can visit this to get an idea on hw to convert impropr fract into mixed fracs. http://www.mathsisfun.com/improper-fractions.html

OpenStudy (ajprincess):

Does that help @Gigglez?

OpenStudy (anonymous):

no -.-

OpenStudy (ajprincess):

Which part of it do u nt understand @gigglez?

OpenStudy (anonymous):

all of it yo... i suck at this fraction stuff

OpenStudy (ajprincess):

Let me explain it from the beginning by taking a simple example \[\frac{2}{3}\] The number in the upper part, that is 2 is called as numerator and the one in the lower part, that is 3 is called as denominator. In a fraction if the numerator is greater than the denominator then that fraction is called as improper fraction as in the following example. \[\frac{3}{2}\] Following me till this?

OpenStudy (anonymous):

yeah i got you so far...

OpenStudy (ajprincess):

ok good:)

OpenStudy (ajprincess):

Improper fractions can be expressed as mixed fractions. Example for mixed fraction is \[3\frac{2}{3}\] Nw let us see how to convert \[\frac{3}{2}\] into a mixed fraction. To do that we divide 3 by 2. We know that 3 is not divisible by 2. When we divide it we get one as a remainder and 1 as quotient.  write the quotient, in this case 1 followed by the fraction having remainder 1 as numerator and 2 as denominator. That is \[\frac{3}{2}=1\frac{1}{2}\] See if u can get this. If not please do ask:)